# BWTP model for hadrons $p_{\rm T}$-spectra and elliptic flow $v_{2}$ in AA collisions at the LHC

### v1

@@ VI.1 Hadrons $p_{\textrm{T}}$-spectra, integrated yields, and $\langle p_{\textrm{T}}\rangle$
Figure 8 displays the ratios of the $p_{\textrm{T}}$-integrated yields of various hadrons as a function of the charged-particle multiplicity $M_{\rm ch}$ (defined by Eq. (2)) in Pb–Pb collisions. It is well known that such ratios have smooth behavior versus $M_{\rm ch}$ and are almost independent of the collision system and energy [11, 32]. Corresponding ratios in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [32] are not shown in Fig. 8 for its clearity. Our model curves in this case, as well as for OO collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.36 TeV, coincide practically with the ones for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV. The model describes well all the presented ratios. In particular, it describes the observed enhancement of yields of hadrons containing strange (anti)quarks with respect to pion yields when going from peripheral to central collisions. Enhancement is stronger for hadrons containing more strange (anti)quarks, according to Eq. (20). Figure 8 displays also the suppression of ${\mathrm{K}}^{*0}$and $\rho^{0}$ resonance yields when going from peripheral to central collisions. It is the manifestation of the low-$p_{\textrm{T}}$ suppression of these resonances (see Fig. 7) due to the rescatterings of their decay products in the hadronic medium. The $\phi$ resonance does not have such a suppression since it lives longer than this medium [11].
Next, we show in Fig. 9 the fit of the data on charged-particle $\langle p_{\textrm{T}}\rangle$ versus $M_{\rm ch}$ in Pb–Pb and Xe–Xe collisions [33]. These data involve much more values of $M_{\rm ch}$
@@
Figure 9: Fit of charged particles $\langle p_{\textrm{T}}\rangle$ as a function of their multiplicity density in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 and 5.02 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [33]. The dashed line shows our prediction for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
than the $p_{\textrm{T}}$-spectra measurements [37, 63]. Our prediction for OO collisions is shown too. The sharp rise in model curves at very high $M_{\rm ch}$, related to factor $f_{uc}$ of Eq. (IV), is demonstrated in detail in Fig. 10 with the Pb–Pb data [34] at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for the normalaized $\langle p_{\textrm{T}}\rangle$ and $M_{\rm ch}$ to the corresponding values for 0-5% centrality. Our predictions for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 TeV, Xe–Xe at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV, and OO at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV are also shown. Overall, a good description of the data can be seen.
@@ VI.2 Hadrons $p_{\textrm{T}}$-differential and averaged $v_{2}$
It should be noted that the elliptic flow $v_{2}$ data considered in present paper [32, 51, 52, 53, 54, 55, 56, 60, 61, 64, 67] were measured mostly by the two-particle cumulant method and partly by other close methods that give compatible results within the data uncertainties. This method assumes a minimum pseudo-rapidity gap $\Delta\eta$ between the two particles, meant to suppress the non-flow correlations, for which we have chosen the largest available value in any dataset used.
Figure 11: Fit of pion, kaon, proton, and charged-particle $v_{2}$ as a function of $p_{\textrm{T}}$ in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for four centrality intervals, measured by ALICE [52, 53] and CMS [54] experiments. The dashed lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figures 11, 12, and 13 demonstrate our fits of $p_{\textrm{T}}$-differential $v_{2}$ for various hadrons and different centralities in Pb–Pb [51, 52, 53, 54] and Au–Au [67] collisions, as well as some features of such data [11, 60]. A similar good fit is also obtained for the $v_{2}$ data in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [32, 64]. The $v_{2}$ ($p_{\textrm{T}}$) has a wide peak at $p_{\textrm{T}}\sim$ 1–7 GeV/c which height is increasing with the percentage of centrality up to $\sim$50% and then decreasing for more peripheral collisions. This is related to the collision geometry [60] and can be explained in present model by the similar centrality dependence of $\varepsilon_{2}$, defined by Eqs. (7) and (21). The peak is shifted to higher $p_{\textrm{T}}$ for heavier particles. It is the well-known mass-ordering effect due to the radial flow [11, 60]. Such a shift also occurs with increasing centrality (or $M_{\rm ch}$) due to the increasing radial flow velocity $\beta_{\textrm{T}}$ in Eq. (5) caused by the decreasing exponent $k$ defined in Eq. (17). Other known feature of $v_{2}$ is the meson-baryon particle type grouping [11, 60], clearly visible in Fig. 12. This is described in present model by using different parameterizations for mesons and baryons in Eq. (26). Another effect is the increase of $v_{2}$ with $\sqrt{s_{\mathrm{NN}}}$ [11] which can be seen by comparing Figs. 11 and 13. It is mainly related to the increase of $\beta_{\textrm{T}}$ caused by the decreasing $k$ in Eq. (17) with increasing $\sqrt{s_{\mathrm{NN}}}$ (or $M_{\rm ch}$). For the fit shown in Fig. 13 we have used the $\pi^{-}$, $\mathrm{K}^{-}$, and $\mathrm{\overline{p}}$ dataset since it includes more
@@
data-points than the $\pi^{+}$, $\mathrm{K}^{+}$, and $\mathrm{p}$ dataset [67]. Note also that the dashed lines in Fig. 11 show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
It is important to note that the BW term of our model (see Eq. (II)) gives, for the most of considered hadrons, the main contribution to $v_{2}$ in the lower part of its whole $p_{\textrm{T}}$-region up to the center of the peak. The Ts and Po terms, with the use of Eq. (25), describe $v_{2}$ in the rest of this region. However, for hadrons containing heavy quarks, for which the Ts term is zero and the BW term is suppressed by a factor $\sim\exp(-m_{\textrm{T}}/T)$, the Po contribution into $v_{2}$ dominates everywhere (like for $p_{\textrm{T}}$-spectra [1]).
@@
Finally, let us discuss the $p_{\textrm{T}}$-averaged $v_{2}$, defined via Eq. (1) with integration by $p_{\textrm{T}}$ in the numerator and denominator. Figure 14 shows our fit for the charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of $M_{\rm ch}$ [61] (left panel) and centrality fraction $x_{c}$ (defined in Sec. III[53] (right panel). Although both datasets were obtained by the two-particle cumulant method, they are independent measurements since Refs. [61] and [53] use $\Delta\eta>1.4$ and $\Delta\eta>1$, respectively. Very good fits of the Pb–Pb and Xe–Xe data can be seen in Fig. 14. The black lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
Figure 14: Fit of charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of multiplicity density [61] (left panel) and $x_{c}$ [53] (right panel) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 and 2.76 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV. The black lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. The dashed lines correspond to $v_{2}$ averaged in the $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c.
@@
For comparison, dashed lines are shown, corresponding to $v_{2}$ averaged in the much larger $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c. The left panel of Fig. 14 shows that our model describes well the different $M_{\rm ch}$ (or centrality) dependence of $v_{2}$ in Xe–Xe and Pb–Pb collisions. This is due to the presence of the parameter $f_{d}$, related to the non-spherical form of 129Xe nucleus, and of the A-dependence in Eq. (21). Also important is the smaller $\beta_{\textrm{T}}$ (caused by the larger $k$ from Eq. (17)) for collisions of smaller nuclei. These A-dependence and smaller $\beta_{\textrm{T}}$ are the main reasons for the lower values and flatter centrality (or $M_{\rm ch}$) dependence of the predicted $v_{2}$ in OO collisions with respect to Pb–Pb collisions, shown in Figs. 11 and 14.
@@ VII Conclusions
Thus, a more general and parameterized version of BWTPM [1] is presented, which has much fewer free parameters, but successfully describes not only the midrapidity $p_{\textrm{T}}$-spectra, but also elliptic flow $v_{2}$ of various hadrons with arbitrary $p_{\textrm{T}}$ in AA collisions for any centrality at the LHC energies. It is shown that the model also works at much lower RHIC energies. For all considered particles, from pions to charmonia, BWTPM allows us to calculate easily the midrapidity $p_{\textrm{T}}$-spectra and $v_{2}$ for new AA collisions with any A and centrality at high energies. Predictions are given for the OO collisions carried out recently at the LHC. To calculate the $p_{\textrm{T}}$-spectra or $v_{2}$ for new particles, not considered here, one needs some data (e.g., $p_{\textrm{T}}$-spectrum for one centrality or $\langle p_{\textrm{T}}\rangle$ versus centrality) to fix at least two particle-type dependent parameters $f_{BW}$ and $f_{2}$.
BWTPM can be efficiently used by experimenters to quickly check new data on hadron $p_{\textrm{T}}$-spectra and $v_{2}$, and by theorists to test their more general and sophisticated theories and models of hadron production in AA collisions. Similarly, the simple parameterization for the charged-particle multiplicity density at midrapidity in AA or pp collisions presented in Sec. III can also be used.
@@ Acknowledgments
I thank Jurgen Schukraft for valuable comments on the BWTPM parameterization.
@@
The data supporting the findings of this article are publicly available and cited in the bibliography.

### v3

@@ VI.1 Hadrons $p_{\textrm{T}}$-spectra, integrated yields, and $\langle p_{\textrm{T}}\rangle$
Figure 8 displays the ratios of the $p_{\textrm{T}}$-integrated yields of various hadrons as a function of the charged-particle multiplicity $M_{\rm ch}$ (defined by Eq. (2)) in Pb–Pb collisions. It is well known that such ratios have smooth behavior versus $M_{\rm ch}$ and are almost independent of the collision system and energy [11, 33]. Corresponding ratios in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [33] are not shown in Fig. 8 for clarity. The model curves in this case, as well as for OO collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.36 TeV, coincide practically with the ones for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV. The model describes well all the presented ratios. In particular, it describes the observed enhancement of yields of hadrons containing strange (anti)quarks with respect to pion yields when going from peripheral to central collisions. Enhancement is stronger for hadrons containing more strange (anti)quarks, according to Eq. (20). Figure 8 displays also the suppression of ${\mathrm{K}}^{*0}$ and $\rho^{0}$ resonance yields when going from peripheral to central collisions. It is the manifestation of the low-$p_{\textrm{T}}$ suppression of these resonances (see Fig. 7) due to the rescatterings of their decay products in the hadronic medium. The $\phi$ resonance does not have such a suppression since it lives
@@
Figure 9: Fit of charged particles $\langle p_{\textrm{T}}\rangle$ as a function of their multiplicity density in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 and 5.02 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [34]. Dashed line shows the prediction for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
longer than this medium [11].
Next, Fig. 9 shows the fit of the data on charged-particle $\langle p_{\textrm{T}}\rangle$ versus $M_{\rm ch}$ in Pb–Pb and Xe–Xe collisions [34]. These data involve much more values of $M_{\rm ch}$ than the $p_{\textrm{T}}$-spectra measurements [38, 64]. The prediction for OO collisions is shown too. The sharp rise in model curves at very high $M_{\rm ch}$, related to factor $f_{uc}$ of Eq. (IV), is demonstrated in detail in Fig. 10 with the Pb–Pb data [35] at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for the normalaized $\langle p_{\textrm{T}}\rangle$ and $M_{\rm ch}$ to the corresponding values for 0-5% centrality. The predictions for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 TeV, Xe–Xe at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV, and OO at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV are also shown. Overall, a good description of the data can be seen.
@@ VI.2 Hadrons $p_{\textrm{T}}$-differential and averaged $v_{2}$
It should be noted that the elliptic flow $v_{2}$ data considered in present paper [33, 52, 53, 54, 55, 56, 57, 61, 62, 65, 68] were measured mostly by the two-particle cumulant method and partly by other similar methods that give compatible results within the data uncertainties. This method assumes a minimum pseudo-rapidity gap $\Delta\eta$ between the two particles, meant to suppress the non-flow correlations, for which the largest available value in any dataset used was chosen.
Figure 11: Fit of pion, kaon, proton, and charged-particle $v_{2}$ as a function of $p_{\textrm{T}}$ in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for four centrality intervals, measured by ALICE [53, 54] and CMS [55] experiments. The dashed lines show the predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figures 11, 12, and 13 demonstrate the fits of $p_{\textrm{T}}$-differential $v_{2}$ for various hadrons and different centralities in Pb–Pb [52, 53, 54, 55] and Au–Au [68] collisions, as well as some features of such data [11, 61]. A similarly good fit is also obtained for the $v_{2}$ data in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [33, 65]. The $v_{2}$ ($p_{\textrm{T}}$) has a wide peak at $p_{\textrm{T}}\approx$ 1–7 GeV/c whose height increases with the percentage of centrality up to $\sim$50% and then decreases for more peripheral collisions. This is related to the collision geometry [61] and can be explained in the present model by the similar centrality dependence of $\varepsilon_{2}$, defined by Eqs. (7) and (21). The peak is shifted to higher $p_{\textrm{T}}$ for heavier particles. It is the well-known mass-ordering effect due to the radial flow [11, 61]. Such a shift also occurs with increasing centrality (or $M_{\rm ch}$) due to the increasing radial flow velocity $\beta_{\textrm{T}}$ in Eq. (5) caused by the decreasing exponent $k$ defined in Eq. (17). Other known feature of $v_{2}$ is the meson-baryon particle type grouping [11, 61], clearly visible in Fig. 12. This is described in present model by using different parametrizations for mesons and baryons in Eq. (26). Another effect is the increase of $v_{2}$ with $\sqrt{s_{\mathrm{NN}}}$ [11], which can be seen by comparing Figs. 11 and 13. It is mainly related to the increase of $\beta_{\textrm{T}}$ caused by the decreasing $k$ in Eq. (17) with increasing $\sqrt{s_{\mathrm{NN}}}$ (or $M_{\rm ch}$). For the fit shown in Fig. 13 the $\pi^{-}$, $\mathrm{K}^{-}$, and $\mathrm{\overline{p}}$ dataset was used since it includes more data
@@
points than the $\pi^{+}$, $\mathrm{K}^{+}$, and $\mathrm{p}$ dataset [68]. Note also that the dashed lines in Fig. 11 show the predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
It is important to note that the BW term of the present model (see Eq. (II)) gives, for most of considered hadrons, the main contribution to $v_{2}$ in the lower part of its whole $p_{\textrm{T}}$-region up to the center of the peak. The Ts and Po terms, with the use of Eq. (25), describe $v_{2}$ in the rest of this region. However, for hadrons containing heavy quarks, for which the Ts term is zero and the BW term is suppressed by a factor $\sim\exp(-m_{\textrm{T}}/T)$, the Po contribution to $v_{2}$ dominates everywhere (as for $p_{\textrm{T}}$-spectra [1]).
@@
Finally, let us discuss the $p_{\textrm{T}}$-averaged $v_{2}$, defined via Eq. (1) with integration by $p_{\textrm{T}}$ in the numerator and denominator. Figure 14 shows the fit for the charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of $M_{\rm ch}$ [62] (left panel) and centrality fraction $x_{c}$ (defined in Sec. III[54] (right panel). Although both datasets were obtained by the two-particle cumulant method, they are independent measurements since Refs. [62] and [54] use $\Delta\eta>1.4$ and $\Delta\eta>1$, respectively. Very good fits of the Pb–Pb and Xe–Xe data can be seen in Fig. 14. The black lines show the predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
@@
Figure 14: Fit of charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of multiplicity density [62] (left panel) and $x_{c}$ [54] (right panel) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 and 2.76 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV. The black lines show the predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. The dashed lines correspond to $v_{2}$ averaged in the $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c.
@@
For comparison, dashed lines are shown, corresponding to $v_{2}$ averaged in the much larger $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c. The left panel of Fig. 14 shows that the model describes well the different $M_{\rm ch}$ (or centrality) dependences of $v_{2}$ in Xe–Xe and Pb–Pb collisions. This is due to the presence of the parameter $f_{d}$, related to the non-spherical form of 129Xe nucleus, and of the A-dependence in Eq. (21). Also important is the smaller $\beta_{\textrm{T}}$ (caused by the larger $k$ from Eq. (17)) for collisions of smaller nuclei. These A-dependence and smaller $\beta_{\textrm{T}}$ are the main reasons for the lower values and flatter centrality (or $M_{\rm ch}$) dependence of the predicted $v_{2}$ in OO collisions with respect to Pb–Pb collisions, shown in Figs. 11 and 14.
VI.3 Study of hadrons not used in the global fit
To demonstrate the ability of BWTPM to describe the data for other hadrons, not used in the global fit, here I consider the ALICE measurements for the $p_{\textrm{T}}$-spectra of baryon resonances $\Sigma(1385)^{+}$ and $\Sigma(1385)^{-}$ [72] and $\Lambda(1520)$ [73] in Pb–Pb collisions (the sum of the particle and antiparticle yields was measured in each case). The data show that the yields of these short-lived resonances are suppressed in central collisions. This suppression can be quantified via the parameter $f_{r}$ from Eq. (16), as for the ${\mathrm{K}}^{*0}$ and $\rho^{0}$ resonances discussed above. Since the data  [72, 73] on baryon resonances are scarce, especially at low $p_{\textrm{T}}$, one can use for their parameter $f_{r}$ the value given in Table II. Then, to fit these data, there are left
Figure 15: Fit of $p_{\textrm{T}}$-spectra of $\Sigma(1385)^{\pm}$ [72] (left panel) and $\Lambda(1520)$ [73] (right panel) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 and 2.76 TeV, respectively, for three centrality classes.
only three particle-specific free parameters: $f_{BW},f_{2}$, and $h_{n}$. The combined fit of the data [72], shown in the left panel of Fig. 15, gives $\chi^{2}/NDF=6.8/51,f_{BW}=6.14,f_{2}=0.49$, and $h_{n}=1.25$. Note that the value of $h_{n}$ is close to the one given in Table I for hyperons. The $\Lambda(1520)$ data [73] are very scarce at high $p_{\textrm{T}}$, so the fitting function is almost insensitive to the parameter $h_{n}$. The fit of these data, using fixed $h_{n}=1.25$ and two free parameters, gives $\chi^{2}/NDF=5.4/17,f_{BW}=12.8,f_{2}=0.20$, and is displayed in the right panel of Fig. 15.
So, by fitting only few particle-specific parameters, BWTPM allows us to describe the data for new (not considered here) particles produced in AA collisions at different centralities and energies.
@@ VII Conclusions
Thus, a more general and parametrized version of BWTPM [1] is presented, which has much fewer free parameters, but successfully describes not only the midrapidity $p_{\textrm{T}}$-spectra, but also elliptic flow $v_{2}$ of various hadrons with arbitrary $p_{\textrm{T}}$ in AA collisions for any centrality at the LHC energies. It is shown that the model also works at much lower RHIC energies. For all considered particles, from pions to charmonia, BWTPM allows us to easily calculate the midrapidity $p_{\textrm{T}}$-spectra and $v_{2}$ for new AA collisions with any A and centrality at high energies. Predictions are given for the OO collisions carried out recently at the LHC. However, for collisions of isobars and other deformed nuclei, shape-specific parameters should be introduced, such as the $f_{d}$ parameter used above for Xe–Xe collisions. To calculate the $p_{\textrm{T}}$-spectra or $v_{2}$ for new particles, not considered in this study, one needs some data (e.g., $p_{\textrm{T}}$-spectrum for at least one centrality or $\langle p_{\textrm{T}}\rangle$ versus centrality) to fix the particle-specific parameters, as shown for the strange baryon resonances in Sec. VI.3.
In the future, it is planned to employ BWTPM to describe the available data on the production of light nuclei (such as deuteron and helion) in AA collisions.
BWTPM can be efficiently used by experimenters to quickly check new data on hadron $p_{\textrm{T}}$-spectra and $v_{2}$. Similarly, the simple parametrization for the charged-particle multiplicity density at midrapidity in AA or pp collisions presented in Sec. III can also be used.
@@ Acknowledgments
I thank Jurgen Schukraft for valuable comments on the BWTPM parametrization. I also thank the anonymous referee for important comments that led to further discussion of deformed nucleus collisions and BWTPM parameters as well as the addition of Sec. VI.3.
@@
The data supporting the findings of this article are openly available [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 21, 27, 34, 35, 37, 38, 41, 42, 43, 39, 40, 44, 45, 46, 49, 50, 47, 48, 51, 52, 55, 56, 53, 54, 57, 58, 59, 60, 61, 62, 33, 63, 64, 65, 66, 67, 68, 69, 70, 72, 73].